The average length of time for students to register for classes at a certain college has been 50 minutes.A new registration procedure using modern computing machines is being tried.If a random sample of 12 students had an average registration time of 42 minutes with a standard deviation of 11.9 minutes under new system , test the hypothesis that the population mean is now less than 50 using 0.01 level of significance.
1
Expert's answer
2020-03-21T14:01:15-0400

Since the population standard deviation is unknown, the mean follows a t(n-1) distribution.

"H_0 \mu=50"

"H_a \mu<50"

"t-value= \frac{\bar{X}-\mu}{\frac{s}{\sqrt{n}}}"

"= \frac{42-50}{\frac{11.9}{\sqrt{12}}}=-2.32881"

The critical region (left sided)

Obtained from =T.INV(0.01,11) excel formula.

"t_{(0.01,11)}=-2.7181"

Since absolute t-value (2.33) is less than the absolute critical value (2.72), we fail to reject the null hypothesis and conclude that there is no sufficient evidence to conclude that the mean I less than 50.